OpenStudy (anonymous):

Solve for x -5|x+1|=10

OpenStudy (one098):

First add -5 from both sides. Then subtract 1 from both sides.

OpenStudy (anonymous):

\[ x -5|x+1|=10 \\ 5|x+1 |= x-10\\ |x+1 |=\frac 1 5 ( x-10)\\ x+1=\pm \frac 1 5 ( x-10)\\ \]

OpenStudy (anonymous):

You get \[ x=-\frac {15}4\\ x= \frac 5 6 \]

OpenStudy (anonymous):

An easier way is to solve \[ x-5(x+1)=10 \\ and\\ x+5(x+1)=10\]